If the height of a triangle with a base of 8cm is tripled, its Area is by 96 square centimeters. Find the height of the triangle?
Thank you

Answer by lwsshak3(11628)

If the height of a triangle with a base of 8cm is tripled, its Area is increased by 96 square centimeters. Find the height of the triangle?
***
Area of triangle=1/2 base*height
let h=height of triangle
1/2*8*3h-1/2*8*h=96
12h-4h=96
8h=96
h=12
height of the triangle=12 cm